OpenStudy (chaise):

Because I'm bored. A man is 20km out to sea when his boat brakes down, he has no communication devices and has no choice but to swim/run to get help. Help is 60km down the coast. If the man can swim at 6km/h and can run at 12km/h. How long will it take him to get help?

OpenStudy (chaise):

I should really rephrase it - What is the minimum amount of time it will take for him to get help, and what direction from north should he swim in to get help, to achieve the shortest possible time? 

OpenStudy (saifoo.khan):

10 hours?

OpenStudy (mimi_x3):

LOL, if your bored please post an easier question that isn't problem solving ?

OpenStudy (saifoo.khan):

Huang drew this.

OpenStudy (mimi_x3):

LOL, I can draw better than that hehe

OpenStudy (chaise):

This question in my opinion is very easy :) Saifoo could do it

OpenStudy (mimi_x3):

He's a fail legend ofc he can't do it xD

OpenStudy (saifoo.khan):

is the answer 10?? mini is just jealous.

OpenStudy (akshay_budhkar):

ME~NI LOL

OpenStudy (akshay_budhkar):

to reach there in shortest possible time he musst swim straight

OpenStudy (mimi_x3):

LOL, Im not MINI, its MIMI, if your not blind. Hell no, Im jealous, jealousy is not in my vocab.

OpenStudy (chaise):

 \[d1 = \sqrt{x^2 + 20^2}\] \[d1 = \sqrt{x^2-400}\] \[d2 = 65-x\] \[t= \frac{d}{s}\] \[t =\frac{d1}{t1}+\frac{d2}{t2}\] \[t = \frac{\6sqrt{x^2-400}}{6}+\frac{65-x}{12}\] \[\frac{dt}{dx}= \frac{x}{(6 \sqrt{x^2-400} }-1/12\] At min/max the derivative =0 \[0=\frac{x}{(6 \sqrt{x^2-400} }-1/12\] \[x = 20\] \[65-x=45\] He must hit the shore 45km from help in order to to help in the fastest time.